A ready-mix concrete company plans to buy a new fleet of 10 delivery trucks at $185,000 each. They are 5 year assets as...

  1. A ready-mix concrete company plans to buy a new fleet of 10 delivery trucks at $185,000 each. They are 5 year assets as per MACRS. The gross (pre-tax) income from operating these trucks is shown below. Fill in the rest of the table to determine their post-tax income flow. The company is in a 25% tax bracket.

    Year

    Pre-Tax Income $

    Depreciation $

    Taxable income $

    Tax $

    Post-Tax Income $

    1

    400,000

    2

    430,000

    3

    450,000

    4

    50,000*

    5

    300,000

    6

    250,000

    * A bad year for the ready-mix business
Answer #1

With 5-Year MACRS, the depreciation rates in years 1 to 6 are 20%, 32%, 19.2%, 11.52%, 11.52%, and 5.76%.

Depreciation amount in each year = price per truck * number of trucks * depreciation rate.

1 2 3 4 5 6 B Pre-tax income 400,000 430,000 450,000 50,000 300,000 250,000 C D E F Depreciation Depreciation Taxable Post-ta

AB F 1 Year Pre-tax income Depreciation rate Depreciation $ Taxable income Tax 21 400000 0.2 =185000*10*C2 =B2-D2 =E2*25% 3 2